# HTESP (High-throughput electronic structure package): a Package for the high-throughput ab initio calculations

### v1

@@ Abstract
We introduce HTESP (High-Throughput Electronic Structure Package), an automated tool designed for high-throughput $ab$ $initio$ calculations. HTESP simplifies the entire workflow, including data extraction, input files generation, calculation submission, result collection, and plotting. The package is implemented in Python and Bash languages. In this paper, we provide detailed information about the package, its installation process and some illustrative examples to demonstrate its usage. Additionally, the package includes comprehensive online documentation [https://neraaz.github.io/HTESP/] on input parameters and tutorials to assist users.
@@ I Introduction
High-throughput $ab$ $initio$ calculations of a wide range of properties for ground state and beyond, using density functional theory (DFT) and related methods, are currently receiving significant attention. This heightened interest is the result of active research in data-driven approaches aimed at discovering new materials with desirable properties. The accumulation of extensive theoretical data through computations enables data-driven discovery by constructing machine learning (ML) and artificial intelligence (AI) models to predict novel compounds and forecast their properties. With the introduction of several databases such as materials project (MP) JOHCRD13_MP ; OCJBGCP15 ; Pymatgen13 , Open Quantum Materials Database (OQMD) SKAMW13 ; oqmd15 , and the automatic flow of materials discovery (AFLOW) aflow12 ; aflux17 , extracting existing data on materials and properties can accelerate the next stage materials discovery targeting different properties, such as electron-phonon coupling (EPC) for phonon-mediated superconductivity (SC). However, obtaining data from these databases, generating tailored input files for further calculations, perform such calculations, and analyze new results can be demanding tasks. Currently, several packages like atomate2 atomate2 , AiiDA aiida , and VASPKIT vaspkit exist to facilitate high-throughput frameworks, but often targeting a single database, one electronic structure code, or specific analyses. In this paper, we present a software package named “HTESP” (High-Throughput Electronic Structure Package), which is written in Python and Bash languages. With the advent of our packages, researchers will have more options to integrate various elements into a single platform. Users can enhance the package’s functionality themselves by utilizing simple bash command-line scripts and corresponding Python scripts.
@@
The process of calculating phonon-mediated SC properties is highly intricate with density functional perturbation theory (DFPT) BDDG01 ; D01 . It encompasses a series of steps, starting from the computation of ground-state properties, phonons, EPC, and various subsequent post-processing tasks, all of which, with isotropic Eliashberg’s approximation, are implemented within the Quantum Espresso (QE) codeQE09 ; QE217 . In addition to computing SC properties such as the overall EPC strength ($\lambda$), logarithmic average of phonon frequency ($\omega_{log}$), and critical temperature of superconductivity (Tc), it is imperative to examine the thermodynamic phase stability (convex hull analysis) and band structures (electronic and phonon dispersion, spectral functions, and more properties) of materials. Conducting these analyses is essential to determine the viability of a material to facilitate phonon-mediated SC. Additionally, it’s worth noting that these properties can be further enhanced by applying pressure, substitutional alloying, or doping for the prospective candidates. While there are some utility codes that support high-throughput computational frameworks, there is not a single package capable of performing all the high-throughput tasks mentioned above especially via command-line interface on EPC calculations. Our HTESP package has been developed to fulfill this need and more. In the package, the Python scripts are written for tasks such as generating initial inputs, extracting data, conducting analysis, and plotting, while the Bash scripts are utilized for computational tasks, file processing during intermediate stages, and data extraction from calculations. Although the package was initially written for performing tasks on phonon-mediated SC calculations, it has various other useful functionalities. For example, one can also compute elastic properties automatically via command-line with algorithm implemented in Pymatgen Pymatgen13 ; ong2010thermal ; ong2008li . Using the HTESP package, we have calculated the elastic properties of EuAl4, SrAl4, and BaAl4, in a recently published workWNC24 . This enables us to establish a connection between the instability in the transverse acoustic phonon mode and the shearing modulus, helping to clarify the presence of charge density wave (CDW) in the first two compounds and its absence in the last one WNC24 . Similarly, one of the functionalities of the code is to conduct structural relaxation on structures obtained by distorting the structure according to the phonon eigenmodes, examining systems under hydrostatic pressure, and employing various electronic smearing techniques. In a recent study NCW24 , we leveraged this capability to uncover the low-symmetry structure of Y2C3 through distortion according to the imaginary phonon modes at $\Gamma$ point, followed by an investigation into its EPC properties. Additionally, besides distortion, the application of pressure and the utilization of large smearing can also stabilize phonons, facilitating the study of superconductivity, which can be automatically conducted using this code. Here is the current list of the tasks that this package can accomplish, as a high-throughput workflow.
@@
  • Retrieving and formatting input files from MP, AFLOW, and OQMD databases for QE and VASPvasp1 ; vasp2 calculations.
@@
  • Conducting phonon and thermodynamic calculations using the phonopy packagephono1 ; phono2 .
  • Executing ground-state calculations to construct thermodynamic phase diagrams (ground state convex hulls)ong2010thermal ; ong2008li .
  • Performing Fermi surface calculations utilizing VASP output file “vasprun.xml” and the IFERMIifermi package for visualization and analysis.
  • Computing elastic propertiesDCAJ15 , investigating magnetic ordering HMLP19 , magnetic anisotropy energy (MAE), substitutions, and charge calculations.
@@
  • Generating input files for Wannier90wannier90 , EPW (anisotropic superconductivity)epw1 ; epw2 , and WannierToolswanniertools calculations, with energies windows provided by users for wannierization.
@@
At present, It offers functionalities like monitoring calculations, aborting jobs, and printing command histories, accessible through straightforward commands. In this paper, we provide an overview to the package, outline the installation process, and demonstrate its capabilities with examples. More detailed documentation, which includes complete descriptions of inputs, commands, and examples, is included with the package’s online documentation.
II Package overview
@@
Figure 1 illustrates a procedural workflow employed within the package. The core command at the heart of the package is “mainprogram”, and it can be run with the additional “process” parameter specified as “mainprogram process.” This “process” parameter can take the form of numerical values, texts, or a combination of both, each representing a distinct operation. To gain a comprehensive understanding of these processes, users can refer to the “mainprogram basicinfo” command and follow the instructions provided within it. The workflow starts with extracting materials information from the database using a central input file named “config.json”, which serves as the core input file for the package.
Listing 1: Skeleton of the config.json file
@@
The package includes a default version of config.json file with predefined values. This file contains essential details about the materials to be searched within the database, parameters required for preparing input files, pseudopotential information, and other relevant data. For QE, the required input parameters are stored within the “pwscf_in” subdictionary. As for VASP, alongside “config.json” users have the alternative of employing the “vasp.in” input file to modify the INCAR file acquired from the MPRelaxSet class in pymatgen. Using these files, one can conduct searches and generate input files through a combination of the “search” and “download” processes. Two files are produced: “mpid-list.in” through the “search” process and “mpid.in” via the “download” process. The primary difference between these two files is that “mpid.in” exclusively contains information about compounds for which input has been generated, constituting a subset of the “mpid-list.in” file. The generated input files for QE are stored as “scf-mpid.in” within the folder named “scf_dir”. For VASP, distinct folders named “Rmpid-name” containing VASP input files are created. Here, “mpid” refers to the materials IDs, while “name” denotes the compound’s name, which can be found in “mpid.in” or “mpid-list.in” files. The “input.in” file enables the execution of various processes for different tasks with the following format.
1
3
200 0
mpid-list.in
phband
DFT = QE
The “input.in” file format includes the starting index on the first line, the last index (excluded from calculations) on the second line, the total number of k-points (200) and cutoff index for high-symmetry path (from the last) on the third line, the tracking file “mpid-list.in” for sequential indices, the plot type denoted by “phband” (phonon band), and the type of calculations: ”QE” or ”VASP”. A batch header file, ‘batch.header‘ [Fig. 2] containing the necessary batch commands related to the computer cluster is required to generate job submission scripts. Comprehensive descriptions of input files and commands are available in the manual or online documentation of the package.
III Installation
III.1 Download Software
git clone https://github.com/Neraaz/HTESP.git
Go to HTESP directory,
cd HTESP
III.2 Conda Environment
conda create --name myenv python==3.9.12
source activate myenv
III.3 Install dependencies
pip install -r requirements.txt
Install Phonopy (optional) to perform Phonopy calculations.
III.4 Install HTESP Package
pip install . (pip install dot)
III.5 Check Executable
which mainprogram
Execute “mainprogram basicinfo” to begin.
python setup.py develop
which mainprogram
III.6 After Installation
export PATH="path_to_HTESP/src/bash:\$PATH"
Provide the path to src file
export PYTHONPATH="path_to_HTESP/src:\$PYTHONPATH"
IV Illustrative Examples
IV.1 Printing information about package
Once installed, you can execute “mainprogram basicinfo” to obtain necessary information about the package functionality. Here are different processes and their corresponding operations.

### v3

# HTESP (High-throughput electronic structure package): a Package for high-throughput ab initio calculations

@@ Abstract
High-throughput $ab$ $initio$ calculations are the indispensable parts of data-driven discovery of new materials with desirable properties, as reflected in the establishment of several online material databases. The accumulation of extensive theoretical data through computations enables data-driven discovery by constructing machine learning and artificial intelligence models to predict novel compounds and forecast their properties. Efficient usage and extraction of data from these existing online material databases can accelerate the next stage materials discovery that targets different and more advanced properties, such as electron-phonon coupling for phonon-mediated superconductivity. However, extracting data from these databases, generating tailored input files for different $ab$ $initio$ calculations, performing such calculations, and analyzing new results can be demanding tasks. Here, we introduce a software package named “HTESP” (High-Throughput Electronic Structure Package) written in Python and Bash languages, which automates the entire workflow including data extraction, input file generation, calculation submission, result collection and plotting. Our HTESP will help speed up future computational materials discovery processes.
@@ I Introduction
High-throughput $ab$ $initio$ calculations of a wide range of properties for ground state and beyond, using density functional theory (DFT) and related methods, are currently receiving significant attention. This heightened interest is the result of active research in data-driven approaches aimed at discovering new materials with desirable properties. The accumulation of extensive theoretical data through computations enables data-driven discovery by constructing machine learning (ML) and artificial intelligence (AI) models to predict novel compounds and forecast their properties [1, 2]. With the introduction of several databases such as materials project (MP) [3, 4, 5], Open Quantum Materials Database (OQMD) [6, 7], and the automatic flow of materials discovery (AFLOW) [8, 9], extracting existing data on materials and properties can accelerate the next stage materials discovery targeting different properties, such as electron-phonon coupling (EPC) for phonon-mediated superconductivity (SC). However, obtaining data from these databases, generating tailored input files for further calculations, performing such calculations, and analyzing new results can be demanding tasks. Currently, several packages like atomate2 [10], AiiDA [11], and VASPKIT [12] exist to facilitate high-throughput frameworks, but often targeting a single database, one electronic structure code, or specific analyses. In this paper, we present a software package named “HTESP” (High-Throughput Electronic Structure Package), which is written in Python and Bash languages. With the advent of our packages, researchers will have more options to integrate various elements into a single platform. Users can enhance the package’s functionality themselves by utilizing simple bash command-line scripts and corresponding Python scripts.
@@
The process of calculating phonon-mediated SC properties is highly intricate with density functional perturbation theory (DFPT) [13, 14]. It encompasses a series of steps, starting from the computation of ground-state properties, phonons, EPC, and various subsequent post-processing tasks, all of which, with isotropic Eliashberg’s approximation, are implemented within the Quantum Espresso (QE) code[15, 16]. In addition to computing SC properties such as the overall EPC strength ($\lambda$), logarithmic average of phonon frequency ($\omega_{log}$), and critical temperature of superconductivity (Tc), it is imperative to examine the thermodynamic phase stability (convex hull analysis) and band structures (electronic and phonon dispersion, spectral functions, and more properties) of materials. Conducting these analyses is essential to determine the viability of a material to facilitate phonon-mediated SC. Additionally, it’s worth noting that these properties can be further enhanced by applying pressure, substitutional alloying, or doping for the prospective candidates. While there are some utility codes that support high-throughput computational frameworks, there is not a single package capable of performing all the high-throughput tasks mentioned above especially via command-line interface on EPC calculations. Our HTESP package has been developed to fulfill this need and more. In the package, the Python scripts are written for tasks such as generating initial inputs, extracting data, conducting analysis, and plotting, while the Bash scripts are utilized for computational tasks, file processing during intermediate stages, and data extraction from calculations. Although the package was initially written for performing tasks on phonon-mediated SC calculations, it has various other useful functionalities. For example, one can also compute elastic properties automatically via command-line with algorithm implemented in Pymatgen [5, 17, 18]. Using the HTESP package, we have calculated the elastic properties of EuAl4, SrAl4, and BaAl4, in a recently published work[19]. This enables us to establish a connection between the instability in the transverse acoustic phonon mode and the shearing modulus, helping to clarify the presence of charge density wave (CDW) in the first two compounds and its absence in the last one [19]. Similarly, one of the functionalities of the code is to conduct structural relaxation on structures obtained by distorting the structure according to the phonon eigenmodes, examining systems under hydrostatic pressure, and employing various electronic smearing techniques. In a recent study [20], we leveraged this capability to uncover the low-symmetry structure of Y2C3 through distortion according to the imaginary phonon modes at $\Gamma$ point, followed by an investigation into its EPC properties. Additionally, besides distortion, the application of pressure and the utilization of large smearing can also stabilize phonons, facilitating the study of superconductivity, which can be automatically conducted using this code. Here is the current list of the tasks that this package can accomplish, as a high-throughput workflow.
@@
  • Retrieving and formatting input files from MP, AFLOW, and OQMD databases for QE and VASP[21, 22] calculations.
@@
  • Conducting phonon and thermodynamic calculations using the phonopy package[23, 24].
  • Executing ground-state calculations to construct thermodynamic phase diagrams (ground state convex hulls)[17, 18].
  • Performing Fermi surface calculations utilizing VASP output file “vasprun.xml” and the IFERMI[25] package for visualization and analysis.
  • Computing elastic properties[26], investigating magnetic ordering [27], magnetic anisotropy energy (MAE), substitutions, and charge calculations.
@@
  • Generating input files for Wannier90[28], EPW (anisotropic superconductivity)[29, 30], and WannierTools[31] calculations, with energies windows provided by users for wannierization.
@@
At present, It offers functionalities like monitoring calculations, aborting jobs, and printing command histories, accessible through straightforward commands. In this paper, we provide an overview to the package, outline the installation process, and demonstrate its capabilities with examples. More detailed documentation, which includes complete descriptions of inputs, commands, and examples, is included with the package’s online documentation [https://neraaz.github.io/HTESP/].
II Methods and Materials
@@
Figure 1 illustrates a procedural workflow employed within the package. A complete installation process of the package is presented in Appendix A. The core command at the heart of the package is “mainprogram”, and it can be run with the additional “process” parameter specified as “mainprogram process.” This “process” parameter can take the form of numerical values, texts, or a combination of both, each representing a distinct operation. To gain a comprehensive understanding of these processes, users can refer to the “mainprogram basicinfo” command and follow the instructions provided within it. The workflow starts with extracting materials information from the database using a central input file named “config.json”, which serves as the core input file for the package [Appendix B].
The package includes a default version of config.json file with predefined values. This file contains essential details about the materials to be searched within the database, parameters required for preparing input files such as magnetism, convergence parameters, optimization algorthms, etc, pseudopotential information, and other relevant data. For QE, the required input parameters are stored within the “pwscf_in” subdictionary. As for VASP, alongside “config.json” users have the alternative of employing the “vasp.in” input file to modify the INCAR file acquired from the MPRelaxSet class in pymatgen. Using these files, one can conduct searches and generate input files through a combination of the “search” and “download” processes. Two files are produced: “mpid-list.in” through the “search” process and “mpid.in” via the “download” process. The primary difference between these two files is that “mpid.in” exclusively contains information about compounds for which input has been generated, constituting a subset of the “mpid-list.in” file. The generated input files for QE are stored as “scf-mpid.in” within the folder named “scf_dir”. For VASP, distinct folders named “Rmpid-name” containing VASP input files are created. Here, “mpid” refers to the materials IDs, while “name” denotes the compound’s name, which can be found in “mpid.in” or “mpid-list.in” files. Furthermore, a “download.csv” file is generated within the “download” folder. The columns in this CSV file match the properties listed under the prop keyword [Appendix E]. You can use the MlProcess class in the ml_processing.py script [from ml_processing import MlProcess] to easily prepare basic input files for machine learning studies by writing some python codes. Currently, one can use the class to create composition-based features with Matminer [32], generating an ‘id_prop.csv‘ file that includes materials ID and properties necessary for models like Crystal Graph Convolutional Neural Networks (CGCNN) [33] and the Atomistic Line Graph Neural Network (ALIGNN) [34]. Although the code currently lacks a command-line interface for this task, we plan to add this feature in the near future with more functionalities. The “input.in” file enables the execution of various processes for different tasks [Figure 2.]
The “input.in” file format includes the starting index on the first line, the last index (excluded from calculations) on the second line, the total number of k-points (200) and cutoff index for high-symmetry path (from the last) on the third line, the tracking file “mpid-list.in” for sequential indices, the plot type denoted by “phband” (phonon band), and the type of calculations: “QE” or “VASP”. A batch header file, ‘batch.header‘ [Fig. 2] containing the necessary batch commands related to the computer cluster is required to generate job submission scripts. Comprehensive descriptions of input files and commands are available in the manual or online documentation of the package.
II.1 Illustrative Examples
III Results
III.1 Electron-phonon calculations utilizing QE code
In this section, we provide an example of computing the phonon-mediated SC properties of MgB2 using the isotropic Eliashberg approximation[35] and the McMillan-Allen-Dynes formula[36, 37, 38], as implemented in QE code. One can perform high-throughput calculations for a large number of systems using the “mpid.in” file. Detailed theory of computing phonon-mediated SC properties is presented in Appendix F. A basic “config.json” file for initiating QE calculations from a structure file in “.cif” format appears as the JSON file is shown in Appendix D.
Place config.json, batch.header [Fig. 2], and mp-763.cif (material ID for MgB2 in MP) files inside working directory (workdir). The pseudopotentials[39, 40] are stored in the “pp” folder in the “element.upf” format, where element = Mg or B. To generate the job submission script (mainprogram jobscript), one can utilize the “batch.header” file containing necessary batch commands. With this process, the actual commands are appended to this file based on the “command_list” key. Some “run-command.sh” files are created, for example, run-scf.sh, run-elph.sh, and so on. Now, generate input file from the “mp-763.cif” file containing the structure for MgB2 by executing the command “mainprogram download”. This creates scf-mp-763.in file inside the scf_dir directory. “input.in” and “mpid.in” files are also created. Replace “mpid-list.in” with “mpid.in” in “input.in” as;
Figure 2: Input control files for QE calculations
In this context, “v1” serves solely as an identifier, “mp-763” represents the materials ID, and “Mg1B2” is the name of the compound. With this step, we are ready to perform structure relaxation and further processes. Throughout the examples in this paper, we approximate the exchange-correlation energy using the generalized gradient approximation (GGA) by Perdew-Burke-Ernzerhof (PBE)[41].
We provide a detailed step-by-step tutorial for computing EPC properties in Appendix G, starting from structure relaxation. In Figure 3, we present structural and electronic properties computed from our command-line automated calculations. The crystal structure has been plotted with VESTA[42], while the BZ plot is from the AFLOW[43] web interface (see Figure 3(a)). Fermi surface has been plotted with IFermi package utilizing VASP output file “vasprun.xml” (see Figure 3(b)). In Figure 3(c) (left panel), we show the electronic band structure of MgB2, demonstrating the metallic behavior (also shown by the DOS/pDOS plot in Figure 3(c) (right panel)) and a relatively flat band near the Fermi level ($E_{F}$) along the $\Gamma$-A direction, consistent with the findings in the Materials Project database[3]. Bands near $E_{F}$ are dominated by B’s 2p orbital. After each relaxation, the basic structural information about the materials can be checked using the “mainprogram compound” command. An example of the output is shown in Figure 3(d). Eliashberg’s spectral function (Figure 4(c)) shows that the largest contribution of $\lambda$ is coming from the 15-20 THz phonon frequency range, which is also confirmed by the mode-resolved $\lambda$-projected phonon dispersion (Figure 4(a)) with regions of denser green shades. Additionally, the integrated $\lambda$ for Gaussian smearing of 0.01 Ry is close to 0.73, agreeing with earlier theoretical calculations[44, 45]. Similarly, we calculate $\omega_{log}$ to be 762 K, and the critical temperature of 16.8 K using the McMillan-Allen-Dynes formula.
Finally, to analyze the contribution of different atomic species to phonon eigenmodes, we have summed the atomic displacements of each eigenmode for different species ($|\mathbf{e}^{i}_{\nu\mathbf{q}}|^{2}$) as shown in Equation (1)[20], where $\mathbf{e}^{ij}_{\nu\mathbf{q}}$ are the displacements obtained for atom type $i$ from the eigenvector of mode $\nu$ for the $q$-point $\mathbf{q}$ by diagonalizing the dynamical matrix, and the index $j$ ranges from 1 to $N_{i}$, representing the number of atoms of type $i$.
III.2 Combining queries from different database
process
Database Search Download
MP search download
OQMD oqmd-search oqmd-download
AFLOW aflow-search aflow-download
Combine data-combine
Next, we show the calculations with VASP from preparing input files to submitting calculations, using config.json file as shown in Appendix E. We outline the procedure for setting up VASP calculations, involving the extraction of data for various compositions of yttrium-carbon (Y-C) from three different databases, combining the data, and preparing systems with distinct spacegroups for VASP computations. Specifically, we extract compounds with “Y-C” combinations in chemsys modes (“Y”, “C”, and “Y-C” systems) from the MP, OQMD, and AFLOW databases. We filter the MP database to include all combinations with an energy above the ground-state hull by less than 0.08 eV/atom, thereby encompassing a wider range of Y-C compounds. A k-point mesh density of 0.025 is employed. To generate input files, particularly the INCAR, we utilize the “vasp.in” file [Figure  6 (left panel)] to update the INCAR using the MPRelaxSet class of pymatgen. We incorporate keywords with values to update the INCAR, while removing keys without values from the INCAR. In Table 1, we show the commands to search, extract, and prepare input files from these databases, and combine those data according to composition and spacegroup. Specify “VASP” or “vasp” for VASP input files when setting up the “calc” flag. These commands prepare and store INCAR, POSCAR, POTCAR, and KPOINTS in Rmpid-name/relax/ folders for VASP. The materials IDs and compound names are printed in the “mpid.in” files, as shown in Figure 5. Now, merge this data by grouping the structures according to composition and spacegroup to generate a new set of folders and files. The updated “mpid.in” file is generated [Figure 6(right panel)] along with input files and folders inside the “filtered_inputs” folder. In the updated “mpid.in” file, the OQMD entries have been substituted because of their identical spacegroup found in another database. Subsequent examples will focus exclusively on VASP calculations.
III.3 Vasp calculations on Y-C systems for thermodynamic phase stability
Once we have the input files located in the ‘R{mpid}-{name}/relax‘ folder, which contains the VASP input files, and the corresponding ‘mpid‘ and ‘name‘ stored in the ‘mpid.in‘ file, we can begin the structural relaxation process. Using the ‘input.in‘ file with the setting ‘calc = ”VASP”‘, we initiate the relaxation procedure following steps 1 through 3, similar to how it is done in QE. Continue repeating process 2 (updating POSCAR) and process 3 (resubmitting) until the code stops further calculations, indicating that the systems have fully relaxed within a single ionic relaxation. This process addresses situations where a calculation is either terminated during optimization or relaxed to a local minimum. Additionally, the input parameters can be continuously adjusted to handle different types of errors during the optimization process. The relaxed structure can be stored in CIF format inside the ‘cif‘ folder, and the total energies per ion and the number of ionic iterations can be recorded in ‘econv.csv‘ using the ‘mainprogram e0‘ command, as shown in Figure 7(a). By executing ‘mainprogram pd‘, we can compute convexhull phase diagram of Y-C systems via pymatgen. The command produces ‘convexhull.csv‘ [Figure 7(b)] file and ‘convexhull.pdf‘ plot. The convex hull representing the isothermal phase diagram is crucial for determining the stability of compounds at ambient pressure and temperature.
IV Conclusions
In summary, we present the HTESP package, which streamlines data extraction from multiple database and automates ab initio calculations via a command-line interface. We detail the installation process and showcase the package’s high-throughput capabilities through electron-phonon coupling calculations for MgB2 and thermodynamic phase stability convex hull for Y-C system. In addition to its capabilities in phonon-mediated superconductivity and ground state convex hull phase diagram calculations, the package also offers tutorials embedded within the code showing its various other functionalities. This highlights the potential of the HTESP package as a valuable toolkit for researchers, particularly those work on materials exploration with high-throughput screening. In the future, we plan to enhance the code by adding more functionalities, particularly in plotting, expanding the package to handle 2D materials, and integrating the package with widely used machine learning models such as CGCNN[33] and ALIGNN[34]. Through a combination of bash and Python scripts, our package provides valuable assistance to newcomers in navigating different databases and conducting high-throughput ab initio calculations. In addition to the software package, comprehensive examples and online documentation are available to facilitate the learning process (https://github.com/Neraaz/HTESP).
V Declaration of competing interest
The authors declare no competing interests.
VI Acknowledgements
VII Data availability
The raw data required to reproduce these findings are available to download from https://github.com/Neraaz/HTESP/tree/main/examples. The processed data required to reproduce these findings are available to download from https://github.com/Neraaz/HTESP/tree/main/examples.
Appendix A Installation
A.1 Download Software
git clone https://github.com/Neraaz/HTESP.git
Go to HTESP directory,
cd HTESP
A.2 Conda Environment
conda create --name myenv python==3.9.12
source activate myenv
A.3 Install dependencies
pip install -r requirements.txt
Install Phonopy (optional) to perform Phonopy calculations.
A.4 Install HTESP Package
pip install . (pip install dot)
A.5 Check Executable
which mainprogram
Execute “mainprogram basicinfo” to begin [Appendix C].
python setup.py develop
which mainprogram
A.6 After Installation
export PATH="path_to_HTESP/src/bash:\$PATH"
Provide the path to src file
export PYTHONPATH="path_to_HTESP/src:\$PYTHONPATH"
Appendix B Skeleton of the config.json file
@@ A.6 After Installation
Appendix C Printing information about package
Once installed, you can execute “mainprogram basicinfo” to obtain necessary information about the package functionality. Here, the package shows the different processes and their corresponding operations. These key processes encompass further subprocesses, which can be explored by diving deeper into the functionalities of the package.